OpenStudy (anonymous):

solve: sqrt2cos(x+pi/4)-1=0 for 0

OpenStudy (noelgreco):

Still working on this?

OpenStudy (jhannybean):

are ya?

OpenStudy (arnavguddu):

cos (x+pi/4) = cos (x+45 deg) = cos x cos 45 + sin x sin 45 = 1/(sqrt2) * ( cos x + sin x) So.... ur given eq---- in LHS the sqrt 2 cancels out and u get cos x + sin x - 1 = 0 cos x + sin x = 1 got u this far......

OpenStudy (arnavguddu):

rest will be....square the whole thing.... u will get cos2 x + sin2 x + 2 sin x cos x = 1 ..... rest 2 sin x cos x = 0 gives either sin x = 0 or cos x = 0 both gives the solutions for x